Given that a is an arithmetic sequence, a5=2 and a13=-30 what is a62?
Answer by stanbon(75887)

Given that a is an arithmetic sequence, a5=2 and a13=-30 what is a62?
Equations:
a5:: a + 4d = 2
a13: a + 12d = -30
------
Subtract and solve for "d"::
8d = -32
d = -4
-----
Solve for "a":
a + 4d = 2
a - 16 = 2
a = 18
-------
a62 = a + (61)d = 18 + 61*-4 = 18-244 = -226
-------------------
Cheers,
Stan H.
------------------